Imagine that a fridge has dimensions 1m by 1m by 1.9 m tall and insulated walls of 20.0cm thickness ( single layer K=0.15) . If the room is at 20°C and we set the inside temperature to 4°C then find a) the rate of thermal energy transfer b) the electric energy it uses if it has COP 1.75 c) heat energy created by fridge


1
Expert's answer
2021-02-07T19:08:12-0500

Answer

a) the rate of thermal energy transfer

"Q_t=\frac{AK\Delta T}{L}"

"Q_t=\frac{1\times1 \times1.9 \times(20-4) }{0.2}" =152J/°C


b) the electric energy

"Q=1.75\times152=266J"


c) COP 1.75

So heat energy created by fridge

"Q=1.75\times152=266J"